# Performance Improvement of Deorbitalized Exchange-Correlation Functionals

### v1

@@ I Introduction
The enormous impact of Hohenberg-Kohn-Sham density functional theory (DFT) upon computational study of molecular and materials properties is a consequence of the remarkable cost-accuracy balance of DFT when implemented with modern approximate exchange-correlation (XC) functionals. The cost-accuracy trade-off is determined by the intrinsic limits and transferability of the specific XC approximation used. That compromise clearly differs by problem class. A small set of very large molecules can be treated with an XC approximation far up the Perdew-Schmidt “Jacob’s ladder” complexity hierarchy Perdew and Schmidt (2001). In contrast, the computational costs of high-throughput screening calculations on large data sets of such molecules or ab initio molecular dynamics (AIMD) on counterpart condensed phases compel use of lower-rung approximations. See, for example, the discussion in Sections 2.2.21-23 of Ref. 2.
The SCAN meta generalized-gradient approximation (metaGGA) Sun et al. (2015, 2016) and its variants Furness and Sun (2019); Bartók and Yates (2019); Furness et al. (2020), have proven to be effective compromises for accuracy and efficiency in structural and energetic calculations for both solids and molecules  Zhang et al. (2018); Sun et al. (2016) albeit with some limitations Kovács et al. (2019); Mejía-Rodríguez and Trickey (2019). A design advantage of such meta-GGA functionals is the introduction of inherently nonlocal density information through the use of the non-interacting (Kohn-Sham, KS) kinetic energy density Becke and Edgecombe (1990); Sun et al. (2013),
@@
with orbitals $\phi_{i}$ and occupation numbers $f_{i}$. A resultant design advantage is satisfaction of many more exact constraints than possible with a GGA, particularly for the iso-orbital limit.
@@
Despite the successes, there are difficulties of implementation with these metaGGAs that limit their full exploitation. The explicit orbital-dependence introduced by $\tau_{s}$ causes typical practical application to rely on non-local XC potentials in generalized KS (gKS) equations. Solutions to those can be slow relative to pure KS implementations Mejía-Rodríguez and Trickey (2017). This slowness also originates with the complicated form of the functionals and their potentials, their consequent sensitivity to the density Holzwarth et al. (2022), and concomitant need for dense grids for converged integrals Yao and Kanai (2017); Lehtola and Marques (2022). Calculations sometimes are done non-self-consistently, Zhang et al. (2018); Bonfà et al. (2024) and self-consistent pseudopotentials for these functionals have proven difficult to develop. Holzwarth et al. (2022)
@@
Effort to resolve these issues has led to the redesign of SCAN, first by streamlining performance at the cost of significant constraint noncompliance, Bartók and Yates (2019) then by iso-orbital indicator redefinition, Furness and Sun (2019) and, most recently, in r2SCAN, Furness et al. (2020) by restoration of most of the constraints met by SCAN, while retaining the improvements in numerical performance of prior variants. Nonetheless, some numerical problems remain. Holzwarth et al. (2022); Lehtola and Marques (2022)
@@
Deorbitalization Mejía-Rodríguez and Trickey (2017, 2018, 2020) has been developed as a strategy both to obviate these issues and to retain the interpretive power of the pure KS equation. In deorbitalization, the $\tau_{s}$ dependence used in meta-GGA XC functionals as a local indicator of bond character is replaced with a function of the density $n({\mathbf{r}})$, its gradient $\nabla n$, and Laplacian $\nabla^{2}n$. The result is a pure XC density functional and a pure (local) KS potential.
@@
The obvious deorbitalization requirement is a sufficiently accurate approximate orbital-free kinetic energy density (KED) functional. That is an active area of research. Mi et al. (2023) Practical deorbitalization has come through the methodology of Mejía-Rodríguez and Trickey Mejía-Rodríguez and Trickey (2017, 2018, 2020) (M-RT). They reparametrized existing KED functionals, e.g. Perdew-Constantin (PC) Perdew and Constantin (2007) and others from orbital-free DFT Mi et al. (2023), to match KED-dependent indicator functions of the first 18 isolated atoms. The most generally successful case of M-RT deorbitalization involves the reparametrization (PCopt) of original PC. That gave successful deorbitalizations of both SCAN Mejía-Rodríguez and Trickey (2017, 2018) and r2SCAN Mejía-Rodríguez and Trickey (2020), in the sense of providing what was deemed to be a faithful reproduction of the parent metaGGA error patterns for structural and energetic properties evaluated on standard molecular and crystalline benchmarks.
@@
That faithfulness also has held up in application to materials and outside the original test sets. Rani et al. (2024); Zhang (2024) Other KEDs, e.g. Cancio-Redd (CR)  Cancio and Redd (2017) actually may lead to improved error patterns relative to use of the exact KED for some metaGGAs, but may not for best-in-breed functionals.
@@
A limitation of the M-RT error-pattern criterion of deorbitalization faithfulness is that it can lead to acceptance of deorbitalized XC functionals that deviate noticeably from the exact form in the limit of slowly varying density, i.e., constraint violation. In the context of rethinking the PC functional form, the recent OFR2 study Kaplan and Perdew (2022) emphasized constraint satisfaction, including restoration of correct slowly-varying limit behavior. That recovers fourth-order gradient expansion compliance lost in going from SCAN to r2SCAN Bartók and Yates (2019). The resulting deorbitalization does well in calculations for solids (retaining the numerical advantages of rSCAN) but not as well as r2SCAN-L (r2SCAN deorbitalized by PCopt, the reparametrized version of PC used by M-RTMejía-Rodríguez and Trickey (2017, 2018)) for molecules.
@@
As an aside for clarity, note that a second modification of the original M-RT deorbitalization has been developedFrancisco et al. (2024) to treat metaGGAs that rely on more than one iso-orbital indicator, particularly the Tao Mo functional (TM) Tao and Mo (2016). That is not relevant here.
@@
Alternative strategies to Laplacian-level deorbitalization include use of higher-order spatial derivatives de Silva and Corminboeuf (2014, 2015a, 2015b) and deorbitalization of the exchange potential only Bonfà et al. (2024). Those are of only contextual interest here, as they may have more significant numerical challenges than those arising from the M-RT procedure.
Given success at reproducing the test-case performance of existing orbital-dependent metaGGAs, the challenge is the effective numerical implementation of such a deorbitalized functional. The well-known sensitivity to density Laplacians serves as a caution. See but for one example, the discussion in Ref. 13. Nonetheless, some of the M-RT functionals Mejía-Rodríguez and Trickey (2017, 2018) have shown marked performance improvement over their gKS parents, with time per self-consistent-field (SCF) cycle of the Kohn-Sham solution as much as a factor of three shorter with the VASP code Kresse and Furthmüller (1996); Kresse and Joubert (1999a). But recent work indicates Francisco et al. (2024) that it is quite typical for deorbitalized functionals to require many more self-consistent cycles to achieve SCF convergence than for their parent functionals with gKS, thus offsetting the gains made per cycle.
@@
A usefully accurate description of the KED for deorbitalizing a meta-GGA indicator functional by a simple generalized gradient approximation (GGA) alone does not seem possible Mejía-Rodríguez and Trickey (2017). But, like Laplacian-based KEDFs, Laplacian-based XC deorbitalizers can be ill-behaved with regards to integral convergence and require relatively dense grids. Lehtola and Marques (2022). Moreover, they also can have large, spiky variations in the magnitude of the XC potential. Such variations arise from the high-order spatial derivatives in the X potential that follow from $\nabla^{2}n$Kaplan and Perdew (2022)
@@
The natural, critical question raised by these issues is this: can deorbitalization, in fact, succeed in achieving faster, more reliable metaGGA calculations? At a more detailed level, how does the difficulty of using $\nabla^{2}n$ compare with the difficulties of gKS calculations? For which run conditions, and for which applications, if any, does deorbitalization provide a clear advantage? With respect to computational efficiency, where and how can M-RT deorbitalization be improved?
@@
One clear opportunity is to find ways to reduce numerical instabilities induced by the density-Laplacian. In this work, we take a specific step in that amelioration by intervening in those aspects of the deorbitalized functional that are most problematic, while retaining (or even eventually restoring) important constraints along the way. Our scheme is based on exploitation of the Cancio-Redd Cancio and Redd (2017) and Cancio-Stuart-Kuna Cancio et al. (2016) (CR hereafter) approach of minimizing the structural complexity of KED functionals. We combine that with the constraint satisfaction of the “r2SCAN piece-wise polynomial” (hereafter RPP) deorbitalized indicator function Kaplan and Perdew (2022) used in the OFR2 functional to produce a deorbitalizer with a smoother X potential than OFR2. We show that this functional performs well compared to r2SCAN and to r2SCAN-L (PCopt) and as well or better than OFR2 for solid test cases. It improves significantly on OFR2 for molecular test sets but still is not as good for those as r2SCAN-L(PCopt). We leave until later a more systematic approach to constructing smoothed functionals along with a measure to assess the result.
@@
Secondly, we report a study of timing characteristics, using a single compute node. We show that deorbitalization can be an attractive option for metaGGA calculations, reproducing similar if not better results than original metaGGA for solids and for equation of state calculations, being up to twice as fast in total compute time as gKS calculations. At the same time, we find a number of situations in which the two approaches are about equal in performance and slow compared to GGAs, indicating need for future work.
@@ II.1 MetaGGA Structure
A conventional metaGGA XC functional (ignoring spin-decomposition) is defined by
@@
where both the density $n$ and noninteracting kinetic energy density $\tau_{s}$ are expressed in terms of orbitals $\phi_{i}$ and occupation numbers $f_{i}$, with
@@
In SCAN and some of its descendants, $\tau_{\scriptscriptstyle\rm S}$ enters the XC functional via the iso-orbital indicator $\alpha$,
@@
This may be recognized as the non-negativity of the contribution to the Kohn-Sham kinetic energy derived from Pauli exclusion. Note that the r2SCAN switching function $f_{x}(\bar{\alpha})$ just described is set up to enforce this requirement even if numerical precision errors otherwise would violate it. The indicator $\alpha$ is related trivially to the Pauli kinetic energy density which is the subject of much of orbital-free DFT (OF-DFT)Mi et al. (2023),
@@
To deorbitalize a conventional metaGGA according to M-RT, the exact KS KED is replaced by an approximate orbital-free (OF) semi-local density functional
@@
The deorbitalized metaGGA then is
@@
Comparison of Eqs.  (4) and  (18) illuminates the distinct, key challenges for ordinary versus deorbitalized metaGGA XC numerical implementations. A conventional, orbital-dependent metaGGA introduces a complicated kinetic energy operator in addition to the ordinary local Kohn-Sham potential. A deorbitalized counterpart produces a purely local potential whose construction involves up to fourth spatial derivatives of the density, with concomitant possible numerical difficulties.
@@ II.2 Deorbitalization models
is the dimensionless reduced density Laplacian, partner to the reduced density gradient $p$ defined in Eq. (12). In addition to scaling, other important constraints include $\alpha$ non-negativity [Eq. (13)] and the gradient expansion of $F_{\theta}$ for slowly-varying density. Kirzhnits (1957); Hodges (1973)
KED functionals that have proven particularly useful for deorbitalization share several characteristics. They are proper metaGGAs, in that they depend upon $n$, $p$, and $q$. The design of most has followed, to some degree at least, the strategy employed in the PC KED functional Perdew and Constantin (2007), that is to build a model for the KED in the slowly-varying density limit. Then, because models based on the formally correct gradient expansion in that limit vary much more slowly with $p$ than the iso-orbital limiting case (the von-Weizsäcker functional, $\tau_{\scriptscriptstyle\rm W}$), eventually such models fail the non-negativity constraint Levy and Ou-Yang (1988), $\tau_{s}\geq\tau_{\scriptscriptstyle\rm W}$. That failure is taken to indicate the onset of an iso-orbital or nearly iso-orbital spatial region, and the slowly varying KED is replaced by $\tau_{\scriptscriptstyle\rm W}$ by means of a suitable switching function.
For PC and PCopt (the reparametrized PC used by M-RT Mejía-Rodríguez and Trickey (2017, 2018)), the KED for the slowly-varying limit is a modified fourth-order gradient expansion for which the Pauli enhancement function is
@@
The second-order gradient correction $\Delta F_{\theta}^{(2)}$ consists of a $5p/27$ contribution from the kinetic energy gradient expansion minus a $5p/3$ factor from the von Weizsäcker KED $(\tau_{\scriptscriptstyle\rm W}=(5p/3)\tau_{\scriptscriptstyle\rm TF})$. That yields an overall negative slope with respect to $p$. Together with the possibility of negative $q$, this leads to eventual incipient violation of KED non-negativity which forces the switch from slowly-varying to the iso-orbital model forms.
@@
$F_{\theta}^{\scriptscriptstyle\rm PC}=\alpha_{\scriptscriptstyle\rm PC}=F^{\scriptscriptstyle\rm SV}_{\scriptscriptstyle\rm PC}\,\Theta_{\scriptscriptstyle\rm PC}\left(F^{\scriptscriptstyle\rm SV}_{\scriptscriptstyle\rm PC}\right)$ (28)
The original PC parameter values are $x_{0}=0.5389$, $b=3$. The PCopt values are $x_{0}=1.784720$, $b=0.258304$ Mejía-Rodríguez and Trickey (2017), determined by fitting to the KED $\alpha$ values of small-$Z$ atoms. That reparametrization helps to produce faithful deorbitalization of some metaGGAs but comes at the cost of a somewhat inaccurate form for the slowly-varying limit ($p,q\to 0$).
Some drawbacks of the PC model were uncovered in Ref. Cancio et al., 2016, which provided a modified form to fix them. The switching model $\Theta_{\scriptscriptstyle\rm PC}$ was found to cause unphysical features in the KED in covalent bonds, particularly for systems treated with pseudopotentials. Concurrently, the PC limit for $q\to\infty$ and $p$ finite causes an unphysical treatment of the exponentially decaying density in the asymptotic region of a molecule, characterized by $q\to\infty$ and $p/q\to 1$. The first problem can be resolved by use of a switching factor that obeys $\tau>\tau_{GE}$ and is as smooth as feasible. The second problem can be resolved by recognizing that the second-order gradient expansion for the KED has the scaling behavior of the exact KED for an exponentially decaying density and also is valid for small $p$ and $q$. Thus it can suffice as a bare-bones model for the slowly-varying limit.
Those choices, with further refinement Cancio and Redd (2017), lead to the CR model, namely
$\alpha_{\scriptscriptstyle\rm CR}:=1+\Delta F_{\theta}^{(2)}\Theta_{\scriptscriptstyle\rm CR}\left(\Delta F_{\theta}^{(2)}\right)$ (29)
@@
$H(z)$ is the Heaviside unit step function. The exponent $a=4.0$ produces reasonably close estimates for the total KE of atomsCancio and Redd (2017), while $a=2.0$ produces somewhat smaller Pauli potentials.
The RPP deorbitalizer Kaplan and Perdew (2022) is a PC variant designed specifically for deorbitalizing r2SCAN whilst retaining constraint satisfaction. As such it starts with the same form as Eq. (28),
@@
where $F^{\scriptscriptstyle\rm SV}_{\scriptscriptstyle\rm RPP}$ is, as before, a metaGGA suitable for the slowly varying limit, to wit
@@
This imposes the second-order gradient expansion as the limit for $\to\infty, p/qp/q$ finite, as in CR. Optimal coefficient values were determined against appropriate norms Kaplan and Perdew (2022) to be
@@
The switching functional $\Theta_{\scriptscriptstyle\rm RPP}$ involves the same piece-wise logic as in PC, Eq. (26), but with the nonanalytic switching function $f_{\scriptscriptstyle\rm PC}$, Eq. (27) replaced by the polynomial
@@
Fig. 1 shows how these design choices affect the behavior of the Pauli enhancement factors, i.e., the $\alpha$ to be used in deorbitalization. That figure presents $F_{\theta}$ versus $p$ for the specific value $q=0$. The von-Weizsäcker lower bound value of $F_{\theta}^{\scriptscriptstyle\rm W}=0$ appears as a black solid line. GEA2 denotes the gradient expansion model $1+\Delta F_{\theta}^{(2)}$ that characterizes the slowly varying limit, $p,q\to 0$. The slowly-varying limit for the PC metaGGA starts at the homogeneous electron gas value of 1 at $p=0$ and at first follows along the GEA2 trajectory. But GEA2 transgresses the von-Weizsäcker bound before $p=0.8$. (Recall the negative slope in the $p$-coefficient of the gradient expansion discussed earlier.) The switching function $\Theta_{\scriptscriptstyle\rm PC}$ avoids that crossing by inducing the sudden dip in PC off the GEA2 and subsequent zero value, thus complying with the non-negativity constraint for $F_{\theta}$. That dip is a fairly clear candidate to suspect as the cause of fluctuations in the PC Pauli potential and in the potential of a XC functional deorbitalized with PC. The dip leads to rapid variations in the PC KED in regions of near iso-orbital density, such as covalent bonds, Cancio et al. (2016) which would be exacerbated in calculating the potential. This supposition is shown to be borne out in the following section. In contrast, the CR model (blue-dashed) and SRPP (black dot-dashed) make the transition (from GEA2 to von Weizsäcker bound) in a much smoother fashion, while RPP is an intermediate case. Finally $F_{\theta}$ for PCopt  fails to satisfy the homogeneous electron gas limit ($F\to 1,p,q\to 0$) but does have a reasonably smooth transition.
Figure 1: Pauli enhancement factor $F_{\theta}$ for various KED functionals as a function of reduced density gradient $p$ with reduced Laplacian $q=0$.

### v2

@@ I Introduction
The enormous impact of Hohenberg-Kohn-Sham density functional theory (DFT) upon computational study of molecular and materials properties is a consequence of the remarkable cost-accuracy balance of DFT when implemented with modern approximate exchange-correlation (XC) functionals. The cost-accuracy trade-off is determined by the intrinsic limits and transferability of the specific XC approximation used. That compromise clearly differs by problem class. A small set of very large molecules can be treated with an XC approximation far up the Perdew-Schmidt “Jacob’s ladder” complexity hierarchy [1]. In contrast, the computational costs of high-throughput screening calculations on large data sets of such molecules or ab initio molecular dynamics (AIMD) on counterpart condensed phases compel use of lower-rung approximations. See, for example, the discussion in Sections 2.2.21-23 of Ref. 2.
The SCAN meta generalized-gradient approximation (meta-GGA) [3, 4] and its variants [5, 6, 7], have proven to be effective compromises for accuracy and efficiency in structural and energetic calculations for both solids and molecules  [8, 4] albeit with some limitations [9, 10]. A design advantage of such meta-GGA functionals is the introduction of inherently nonlocal density information through the use of the non-interacting (Kohn-Sham, KS) kinetic energy density [11, 12],
@@
with orbitals $\phi_{i}$ and occupation numbers $f_{i}$. [Throughout we consider only non-relativistic and non-spin-polarized ($f_{i}=0,1,2$) systems and use Hartree atomic units unless otherwise noted.] A resultant design advantage is satisfaction of many more exact constraints than possible with a GGA, particularly for the iso-orbital limit.
@@
Despite the successes, there are difficulties of implementation with these meta-GGAs that limit their full exploitation. The explicit orbital-dependence introduced by $\tau_{\scriptscriptstyle\rm S}$ causes typical practical application to rely on non-local XC potentials in generalized KS (gKS) equations. Solutions to those can be slow relative to pure KS implementations [13]. This slowness also originates with the complicated form of the functionals and their potentials, their consequent sensitivity to the density [14], and concomitant need for dense grids for converged integrals [15, 16]. Calculations sometimes are done non-self-consistently, [8, 17]. Projector-augmented wave (PAW) effective potentials for these functionals have proven difficult to develop [14].
@@
Effort to resolve these issues has led to the redesign of SCAN, first by streamlining performance at the cost of significant constraint noncompliance, [6] then by iso-orbital indicator redefinition, [5] and, most recently, in r2SCAN, [7] by restoration of most of the constraints met by SCAN, while retaining the improvements in numerical performance of prior variants. Nonetheless, some numerical problems remain [14, 16].
@@
Deorbitalization [13, 18, 19] has been developed as a strategy both to obviate numerical issues stemming from use of gKS and to retain the interpretive power of the pure KS equation. In deorbitalization, the $\tau_{\scriptscriptstyle\rm S}$ dependence used in meta-GGA XC functionals as a local indicator of bond character is replaced with a function of the density $n({\mathbf{r}})$, its gradient $\nabla n$, and Laplacian $\nabla^{2}n$. The result is a pure XC density functional and a pure (local) KS potential.
@@
The obvious deorbitalization requirement is a sufficiently accurate approximate orbital-free kinetic energy density (KED) functional. That is an active area of research. [20] A usefully accurate description of the KED for deorbitalizing a meta-GGA indicator functional by a simple generalized gradient approximation (GGA) alone does not seem possible [13]. Inclusion of the density Laplacian proves essential for accurate orbital-free description of the KED at a local level, as demonstrated with the KED density functional provided by Perdew and Constantin [21] (and also Ref. 22). Practical deorbitalization has come through the methodology of Mejía-Rodríguez and Trickey [13, 18, 19] (M-RT). They reparametrized existing Laplacian-level KED functionals, e.g. Perdew-Constantin (PC) [21] and others from orbital-free DFT [20], to match KED-dependent indicator functions of the first 18 isolated atoms. The most generally successful case of M-RT deorbitalization involves the reparametrization (PCopt) of original PC. That gave successful deorbitalizations of both SCAN [13, 18] and r2SCAN [19], in the sense of providing what was deemed to be a faithful reproduction of the parent meta-GGA error patterns for structural and energetic properties evaluated on standard molecular and crystalline benchmarks. Such faithfulness in reproduction defines a deorbitalization goal, one which we follow.
@@
That faithfulness also has held up in application to materials and outside the original test sets. [23, 24] Other KEDs, e.g. Cancio-Redd (CR)  [25] have led, for some meta-GGAs, [26] to better error patterns than those produced with the exact KED, but that has not been seen for best-in-breed meta-GGAs.
@@
A limitation of the M-RT error-pattern criterion of deorbitalization faithfulness is that it can allow acceptance of deorbitalized XC functionals that deviate noticeably from the exact form in the limit of slowly varying density, i.e., constraint violation. In the context of rethinking the PC functional form, the recent OFR2 study [27] emphasized constraint satisfaction, including restoration of correct slowly-varying limit behavior. That recovers fourth-order gradient expansion compliance lost in going from SCAN to r2SCAN [6]. The resulting deorbitalization does well in calculations for solids (retaining the numerical advantages of rSCAN) but not as well as r2SCAN-L (r2SCAN deorbitalized by PCopt, the reparametrized version of PC used by M-RT[13, 18]) for molecules.
@@
As an aside for clarity, note that a second modification of the original M-RT deorbitalization has been developed [28] to treat meta-GGAs that rely on more than one iso-orbital indicator, particularly the Tao Mo functional (TM) [29]. That is not relevant here.
@@
Alternative strategies to Laplacian-level deorbitalization include use of higher-order spatial derivatives [30, 31, 32] and deorbitalization of the exchange potential only [17]. Those are of only contextual interest here, as they may have more significant numerical challenges than those arising from the M-RT procedure.
@@
Given success at reproducing the test-case performance of existing orbital-dependent meta-GGAs, the challenge is the effective numerical implementation of such a deorbitalized functional. The well-known sensitivity to density Laplacians serves as a caution. See for but one example, the discussion in Ref. 13. Laplacian-based XC deorbitalizers, like Laplacian-based KEDFs, can be ill-behaved with regards to integral convergence and require relatively dense grids. [16] Moreover, they also can have large, spiky variations in the magnitude of the XC potential. Such variations arise from the high-order spatial derivatives in the X potential that follow from $\nabla^{2}n$[27]
Nonetheless, some of the M-RT functionals [13, 18] have shown marked performance improvement over their gKS parents, with time per self-consistent-field (SCF) cycle of the Kohn-Sham solution as much as a factor of three shorter with the VASP code [33, 34]. But recent work indicates [28] that it is quite typical for deorbitalized functionals to require many more self-consistent cycles to achieve SCF convergence than for their parent functionals with gKS, thus offsetting the gains made per cycle.
@@
The natural, critical question raised by these issues is this: can deorbitalization, in fact, succeed in achieving faster, more reliable meta-GGA calculations? At a more detailed level, how does the difficulty of using $\nabla^{2}n$ compare with the difficulties of gKS calculations? For which run conditions, and for which applications, if any, does deorbitalization provide a clear advantage? With respect to computational efficiency, where and how can M-RT deorbitalization be improved?
@@
One clear opportunity is to find ways to reduce numerical instabilities induced by the density-Laplacian. In this work, we take a specific step in that amelioration by intervening in those aspects of the deorbitalized functional that are most problematic, while retaining (or even eventually restoring) important constraints along the way. Our scheme is based on exploitation of the Cancio-Redd [25] and Cancio-Stuart-Kuna [22] (CR hereafter) approach of minimizing the structural complexity of KED functionals. We combine that with the constraint satisfaction of the “r2SCAN piecewise polynomial” (hereafter RPP) deorbitalized indicator function [27] used in the OFR2 functional to produce a deorbitalizer with a smoother X potential than OFR2. We show that this functional performs well compared to r2SCAN and to r2SCAN-L (PCopt) and as well or better than OFR2 for solid test cases. It improves significantly on OFR2 for molecular test sets but still is not as good for those as r2SCAN-L(PCopt). We leave until later a more systematic approach to constructing smoothed functionals along with a measure to assess the result.
@@
Secondly, we report a study of timing characteristics, using a single compute node. We show that deorbitalization can be an attractive option for meta-GGA calculations, reproducing similar if not better results than original meta-GGA for solids and for equation of state calculations, being up to twice as fast in total compute time as gKS calculations. At the same time, we find a number of situations in which the two approaches are about equal in performance and slow compared to GGAs, indicating need for future work.
@@ II.1 MetaGGA Structure
A conventional meta-GGA XC functional (ignoring spin-decomposition) is defined by
@@
where both the density $n$ and noninteracting kinetic energy density $\tau_{\scriptscriptstyle\rm S}$ are expressed in terms of orbitals $\phi_{i}$ and occupation numbers $f_{i}$, with
@@
In SCAN and descendants such as r2SCAN, $\tau_{\scriptscriptstyle\rm S}$ enters the XC functional via the iso-orbital indicator $\alpha$,
@@
This may be recognized as the non-negativity of the contribution to the Kohn-Sham kinetic energy derived from Pauli exclusion. Note that the r2SCAN switching function $f_{x}(\bar{\alpha})$ just described is set up to enforce this requirement even if numerical precision errors otherwise would violate it.
The indicator $\alpha$ is related trivially to the Pauli kinetic energy density which is the subject of much of orbital-free DFT (OF-DFT)[20],
@@
To deorbitalize a conventional meta-GGA according to M-RT, the exact KS KED is replaced by an approximate orbital-free (OF) semi-local density functional
@@
The deorbitalized meta-GGA then is
@@
Comparison of Eqs. (4) and (18) illuminates the distinct, key challenges for ordinary versus deorbitalized meta-GGA XC numerical implementations. A conventional, orbital-dependent meta-GGA introduces a complicated kinetic energy operator in addition to the ordinary local Kohn-Sham potential. A deorbitalized counterpart produces a purely local potential whose construction involves up to fourth spatial derivatives of the density, with concomitant possible numerical difficulties.
@@ II.2 Deorbitalization models
is the dimensionless reduced density Laplacian, partner to the reduced density gradient $p$ defined in Eq. (12). Other important constraints include $\alpha$ non-negativity [Eq. (13)] and the gradient expansion of $F_{\theta}$ for slowly-varying density [35, 36].
KED functionals that have proven particularly useful for deorbitalization share several characteristics. They are proper meta-GGAs, in that they depend upon $n$, $p$, and $q$. The design of most has followed, to some degree at least, the strategy employed in the PC KED functional [21], that is to build a model for the KED in the slowly-varying density limit. Then, because models based on the formally correct gradient expansion in that limit vary much more slowly with $p$ than the iso-orbital limiting case (the von-Weizsäcker functional, $\tau_{\scriptscriptstyle\rm W}$), eventually such models fail the non-negativity constraint [37], $\tau_{\scriptscriptstyle\rm S}\geq\tau_{\scriptscriptstyle\rm W}$. That failure is taken to indicate the onset of an iso-orbital or nearly iso-orbital spatial region, and the slowly varying KED is replaced by $\tau_{\scriptscriptstyle\rm W}$ by means of a suitable switching function.
For PC and PCopt (the reparametrized PC used by M-RT [13, 18]), the KED for the slowly-varying limit is a modified fourth-order gradient expansion for which the Pauli enhancement function is
@@
The second-order gradient correction $\Delta F_{\theta}^{(2)}$ consists of a $5p/27$ contribution from the kinetic energy gradient expansion minus a $5p/3$ factor from the von Weizsäcker KED $(\tau_{\scriptscriptstyle\rm W}=(5p/3)\tau_{\scriptscriptstyle\rm TF})$. That yields an overall negative slope with respect to $p$. Together with the possibility of negative $q$, this leads to eventual incipient violation of KED non-negativity which forces the switch from slowly-varying to the iso-orbital model form.
@@
$\alpha_{\scriptscriptstyle\rm PC}=F_{\theta}^{\scriptscriptstyle\rm PC}=F^{\scriptscriptstyle\rm SV}_{\scriptscriptstyle\rm PC}\,\Theta_{\scriptscriptstyle\rm PC}\left(F^{\scriptscriptstyle\rm SV}_{\scriptscriptstyle\rm PC}\right)\;,$ (28)
with the reminder that $\alpha$ and the Pauli enhancement factor $F_{\theta}$ are identical [Eq. (15)]. The original PC parameter values are $x_{0}=0.5389$, $b=3$. The PCopt values are $x_{0}=1.784720$, $b=0.258304$ [13], determined by fitting to the KED $\alpha$ values of small-$Z$ atoms. That reparametrization helps to produce faithful deorbitalization of some meta-GGAs but comes at the cost of a somewhat inaccurate form for the slowly-varying limit ($p,q\to 0$).
Some drawbacks of the PC model were uncovered in Ref. 22, which provided a modified form to fix them. The switching model $\Theta_{\scriptscriptstyle\rm PC}$ was found to cause unphysical features in the KED in covalent bonds, particularly for systems treated with pseudopotentials. Concurrently, the PC limit for $q\to\infty$ and $p$ finite causes an unphysical treatment of the exponentially decaying density in the asymptotic region of a molecule, characterized by $q\to\infty$ and $p/q\to 1$. The first problem can be resolved by use of a switching factor that obeys $\tau>\tau_{GE}$ and is as smooth as feasible. The second problem can be resolved by recognizing that the second-order gradient expansion for the KED has the scaling behavior of the exact KED for an exponentially decaying density and also is valid for small $p$ and $q$. Thus it can suffice as a bare-bones model for the slowly-varying limit.
Those choices, with further refinement [25], lead to the CR model, namely
$\alpha_{\scriptscriptstyle\rm CR}=1+\Delta F_{\theta}^{(2)}\Theta_{\scriptscriptstyle\rm CR}\left(\Delta F_{\theta}^{(2)}\right)$ (29)
@@
$H(z)$ is the Heaviside unit step function. The exponent $a=4.0$ produces reasonably close estimates for the total KE of atoms[25], while $a=2.0$ produces somewhat smaller Pauli potentials.
The RPP deorbitalizer [27] is a PC variant designed specifically for deorbitalizing r2SCAN whilst retaining constraint satisfaction. As such it starts with the same form as Eq. (28),
@@
where $F^{\scriptscriptstyle\rm SV}_{\scriptscriptstyle\rm RPP}$ is, as before, a meta-GGA suitable for the slowly varying limit, to wit
@@
This imposes the second-order gradient expansion as the limit for $\to\infty, p/qp/q$ finite, as in CR. Optimal coefficient values were determined against appropriate norms [27] to be
@@
The switching functional $\Theta_{\scriptscriptstyle\rm RPP}$ involves the same piecewise logic as in PC, Eq. (26), but with the nonanalytic switching function $f_{\scriptscriptstyle\rm PC}$, Eq. (27) replaced by the polynomial
@@
Fig. 1 shows how these design choices affect the behavior of the Pauli enhancement factors, i.e., the $\alpha$ to be used in deorbitalization. That figure presents $F_{\theta}$ versus $p$ for the specific value $q=0$. The von-Weizsäcker lower bound value of $F_{\theta}^{\scriptscriptstyle\rm W}=0$ appears as a black solid line. GEA2 denotes the gradient expansion model $1+\Delta F_{\theta}^{(2)}$ that characterizes the slowly varying limit, $p,q\to 0$. The slowly-varying limit for the PC meta-GGA starts at the homogeneous electron gas value of 1 at $p=0$ and at first follows along the GEA2 trajectory. But GEA2 transgresses the von-Weizsäcker bound before $p=0.8$. (Recall the negative slope in the $p$-coefficient of the gradient expansion discussed earlier.) The switching function $\Theta_{\scriptscriptstyle\rm PC}$ avoids that crossing by inducing the sudden dip in PC off the GEA2 onto the subsequent zero value, thus complying with the non-negativity constraint for $F_{\theta}$. That dip is a fairly clear candidate to suspect as the cause of fluctuations in the PC Pauli potential and in the potential of a XC functional deorbitalized with PC. The dip leads to rapid variations in the PC KED in regions of near iso-orbital density, such as covalent bonds, [22] which would be exacerbated in calculating the potential. This supposition is shown to be borne out in the following section. In contrast, the CR model (blue-dashed) and SRPP (black dot-dashed) make the transition (from GEA2 to von Weizsäcker bound) in a much smoother fashion, while RPP is an intermediate case. Finally $F_{\theta}$ for PCopt  fails to satisfy the homogeneous electron gas limit ($F\to 1,p,q\to 0$) but does have a reasonably smooth transition.
Figure 1: Pauli enhancement factor $F_{\theta}(\equiv\alpha)$ for various KED functionals as a function of reduced density gradient $p$ with reduced Laplacian $q=0$. Models used are: GEA ($1+\Delta F_{\theta}^{(2)}$), PC [Eqs. (23-28)], PCopt(same form, altered coefficients), CR [Eqs. (29-30)], RPP [Eqs. (31-37)] and SRPP [Eq. (38)]. $\alpha^{vW}=0$ by construction.